News by Date
News by Category
Solar System
Stars
White Dwarfs
Supernovas
Neutron Stars
Black Holes
Milky Way Galaxy
Normal Galaxies
Quasars
Galaxy Clusters
Cosmology/Deep Field
Miscellaneous
Press Resources
Status Reports
Press Advisories
Image Releases
Release Guidelines
Image Use Policy
NASA TV
Biographies/Interviews
Web Shortcuts
Chandra Blog
RSS Feed
Chandra Mobile
Chronicle
Email Newsletter
News & Noteworthy
Image Use Policy
Questions & Answers
Glossary of Terms
Download Guide
Get Adobe Reader
Chandra @ NASA
Visit the Chandra pages at the NASA portal
Image Use
Image Use Policy & Request Form
Guidelines for utilizing images, applets, movies, and animations featured in this Web Site.
Getting Hard Copies of Images
Ways to obtain photos, slides, etc of Chandra images.
Operations CXO Status Report

Wednesday 4 August 1999 7.00am EST

The following is an overview of the events that occurred over the last 24 hours. Completed review of plans for IPS #4 with NASA and TRW management. This meeting concluded with confirmation of the Chandra program plan to proceed with a five-minute burn duration using LAE engines #2 & #4.

Completed verification and validation of all commands and software patches required for IPS #4.

Completed Earth Scan 18.

Completed successful dry run of IPS #4.

Measured optical cavity pressure of 8.8 E-06 Torr. Pressure has been reduced from 9.6 E-06 in the last 24 hours

Successfully completed ACIS long form functional test and performed 5 hours of internal calibrations.

Continued HRC calibration activities, but terminated them at the point where optical cavity pressure prohibited further increase in detector operating voltage. This activity will be resumed as pressure continues to decay.

Events scheduled for the next 24 hours

  • Complete IPS #4 with an expected perigee of 5,689 km while retaining the current apogee of 139,108 km (12:35 pm 8/4/99).
  • Continue HRC checkout and ACIS internal source calibration.
  • Develop updated Chandra orbit parameters following IPS #4 and targeting plans for IPS #5.






  • Return to Status Reports